FTS AND ICLAS RECORDS OF PREDISSOCIATED ROVIBRATIONAL STRUCTURE NEAR 10400 CM^{−1} FOR SYMMETRIC ISOTOPOMERS ^{16}O_{3} AND ^{18}O_{3} OF OZONE AND CONTOUR ANALYSIS

\underline{G. WANNOUS}, A.J. BOUVIER, X. CHILLIER, S. CHURASSY, R. BACIS, LASIM, CNRS UMR 5579, Université Claude Bernard-Lyon1, 43 Bd du 11/11/1918, 69622 Villeurbanne Cedex, FRANCE; A.CAMPARGUE, G.WEURAUCH, LSP, CNRS UMR 5588, Université Joseph Fourier de Grenoble, BP 87, 38402 Saint-Martin d’Hères Cedex, FRANCE; R.H. JUDGE, Departement of Chemistry, University of Parkside, Kenosha, WI 53141-2000, USA.

This work is devoted to the study of low-energy quasi-bound states of the O_{3} molecule. Previous analyses^{a,b,c} allowed to caracterize a few vibrational levels of the first excited state ^{3}A_{2}. The absorption spectrum of ^{3}A_{2} in the region of its band 1_{0}^{0} 2_{0}^{2} 3_{0}^{0} presents unassigned features. The theoretical profiles related to the probable connected transitions make difficult an unambigious analysis of the spectrum at room temperature. To decrease the superpositions, absorption spectra at low temperature appeared necessary. Special cooled cells allowed us to record spectra by Fourier Transform Spectroscopy (FTS) and Intra Cavity Laser Absorption Spectroscopy (ICLAS). The predissociation broadening effects are important for these transitions and the rotational structure cannot be observed. A comparison between theoretical and experimental profiles shows two or three types of transitions from the ground state X̃ ^{1}A_{1}. i) One centred at 10360 cm^{−1} corresponds to ^{3}A_{2} ( 0,0,1) ⟵ X̃ ^{1}A_{1} (0,0,0). ii) The others in the 10480 cm^{−1} region could be due to ^{3}B_{2} ⟵ X̃ ^{1}A_{1} and/or ^{3}B_{1} ⟵ X̃ ^{1}A_{1} and, simulations indicate that the upper level of the most intense absorption is probably ^{3}B_{2}. ———————————– aA.J. BOUVIER, G. WANNOUS, S. CHURASSY, R. BACIS, J. BRION, J. MALICET and R.H. JUDGE, Spectrochimica Acta (in press). bA.J. BOUVIER et al. Spectrochimica Acta, \underline{55A}, 2811 (1999) cA.J. BOUVIER et al. J. Mol. Spectrosc., \underline{190}, 189 (1998)